Solid calcium hydroxide, Ca(OH)2, is dissolved in water until the pH of the solution is 10.23. What is the concentration of calcium ion [Ca2+]?
1
Expert's answer
2013-03-01T06:55:55-0500
pH is a measure of the activity of the (solvated) hydrogen ion.pOH is sometimes used as a measure of the concentration of hydroxide ions, OH−, or alkalinity. pOH values are derived from pH measurement.At room temperature pOH ≈ 14 − pH.

pH=14-pOH
pOH=14-pH

pOH=14-10,23=3,77

pOH=-lg[OH-]

[OH]=10^-3,77

The net ionic equation for this process is:

Ca(OH)2 <=> Ca2+ 2 OH-

From this equation you can see that concentration of [OH-] is equal to 2*[Ca2+]

[OH-]=[Ca2+]*2

[Ca2+]=(10^-3,77)/2=8.5*10^-5 M